Consider the graph G of the function f(x) =x^ 2 and the point M(5,2) on the Cartesian plane.

(a)Find the coordinates of a point A on Gand a point B on x-axis, given that M(5,2)is the midpoint of the line segment [AB].

(b) Find the coordinates of a point C on G, given that the distance between M and C is 23


1
Expert's answer
2021-07-28T17:40:34-0400

(a)

"\dfrac{x_A+x_B}{2}=5"

"\dfrac{y_A+y_B}{2}=2"



"x_A+x_B=10"

"x_A^2+0=4"


"x_A=-2, x_B=12"

"A(-2,4), B(12,0)"



"x_A=2, x_B=8"

"A(2,4), B(8,0)"

(b)


"(x_C-5)^2+(y_C-2)^2=23^2"

"y_C=x_C^2"


"x_C^2-10x_C+25+x_C^4-4x_C^2+4=529"


"x_C^4-3x_C^2-10x_C-500=0"


"(x_C-5)(x_C^3+5x_C^2+22x_C+100)=0"


"x_C=5, y_C=25"

"C(5, 25)"



"x_C\approx-4.777, y_C\approx22.820"

"C(-4.777,22.820)"